# Conceptualization, Data Curation

### v2

@@ 4.1 Recurrent Neural Network (RNN) and Long Short-Term Memory (LSTM)
One of the first researches that used LSTM for chemical process fault diagnosis is [130], in this case, no specific feature extraction is needed. Applying batch normalization makes the convergence of the network much faster. To improve the diagnosis using LSTM networks, Han et al. [29] used an iterative method to optimize the number of hidden layer nodes. Kang et al. [38], used RNNs for fault classification. It is shown that RNN can create an even spatial distribution of fault types and classify them accurately, for TEP data a 4-layer RNN could reach the average classification accuracy of 92.5% for all faults, however, it can not detect fault 15. To overcome this problem LSTM network is used and the average accuracy increased to 95%. Compared to the RNN, LSTM has an additional memory cell and three extra gates. To extract long-term dependencies in raw data and detect abnormal values, a Stacked Long Short-Term Memory (LSTM) network with three layers LSTM, is introduced in [122]. As the historical information is stored in the memory, the monitoring system can operate in real time. The average FDR for all TEP faults is 99.49%. Chadha et al. [13] used bidirectional recurrent neural networks for fault diagnosis, enabling the addition of more information about adjacent points into a training sequence than their unidirectional counterparts. They introduced a novel data preprocessing and restructuring for efficient network training. The average FDR for 21 faults of TEP is 88.14%. To overcome nonlinearity, and, the high degree of spatio-temporal correlations, and incomplete characterization of the uncertainty in process noise and dynamics, a novel Bayesian recurrent neural network (BRNN) BRNN-based FDI method is proposed in [83], where we can have a probabilistic model of input-output mapping. This method can be used for fault propagation path analysis as well. To improve the feature extraction of unidirectional RNNs that only consider positive direction, bidirectional RNN (BiRNN) models are used by [124] for FDD of TEP. Simple RNN unit (SRU), LSTM, and GRU are used in making bidirectional networks with the average FDR of 91.0%, 92.3% and 92.7% respectively.
@@
Mirzaei et al. [58], compared LSTM and GRU networks for FDD of chemical processes by using t-SNE visualization. GRU performed better than LSMT, especially for fault 15 which is very hard to detect. The average accuracy of LSTM and GRU for all faults are 95% and 95.9% respectively.
@@
Yuan et al. [117] used GRU to extract the information from correlated sequential data; a moving horizon is used at the input of the GRU to divide the raw data into the sequence units. To overcome the covariate shift and improve the network’s learning efficiency, batch normalization is used. BN-based GRU is faster and more accurate. GRU and BN-GRU accuracy is 83.42% and 87.36%, respectively. An end-to-end model fusion feature learning method (MCNN-DBiGRU) is introduced in [136], a feature-aligned multi-scale extraction model (MCNN) followed by a deep bidirectional mechanism that extracts time series features in the data, allowing the recurrent neural network to process input features both from past to future and from future to past. The experiments on all the fault samples of TEP achieve the precision of 94.11%.
@@ 4.2 Convolutional Neural Network (CNN)
Convolutional Neural Networks (CNNs) were introduced in the late 1980s and initially used to detect patterns in images [39]. These powerful feature extraction networks are widely used for fault detection. A schematic of a CNN is shown in Figure 6.
Figure 6: A schematic of the CNN layers. [3]
@@
One of the first applications of using CNN for TEP fault detection and diagnosis is introduced in [96]. In this approach, a deep convolutional neural network (DCNN) with dropout extracts features from time series data. The implemented method reached the average FDR of 88.2% for 20 faults of TEP, however, this model highly relies on historical fault data samples and is not very effective for super complex chemical processes with thousands of variables. Some modifications to the CNN network are done in [12] to increase the generalization ability of the model, improving feature extraction abilities and reducing the overfitting such as residual connections, inception module, and dilation convolution. With these improvements, the F1 score of hard detectable faults 3,9,13,15 is improved to 0.91 . Frequency domain data contain valuable information that is usually neglected in the traditional FDD of chemical processes. In [120], the Fast Fourier Transform (FFT) is used for TEP feature extraction, and the output data is fed into a convolutional neural network which is used as a binary classifier. The FDR of this model on TEP data considering all faults except 3,9,15,16 is 96.9%. In CNN all the features have the same importance in the classification process. Multi-head attention CNN (MACNN) [20] are introduced to distinguish the importance of different features, where the extracted features from convolution layers are fed into the multi-head attention (MA) layers, finally a softmax classifier is used to classify the different faults, the performance is improved compared to feedforward neural network (FNN) and CNN models. Multiblock temporal convolutional network (MBTCN) for FDD is introduced in [30] using the idea of "local extraction and global integration", it considers the cross-correlation and the temporal-correlation in the multivariate processes’ data. In this method, most related process variables are divided into the same sub-blocks where TCN extract the features. It achieved 96.2% accuracy in TEP considering all fault classes but 3,9,15. In [115], a multichannel one-dimensional convolutional neural network (MC1-DCNN) was used for FDD of the TEP process using extracted features from the Wavelet transform. Wavelet transform is a powerful tool for extracting time and frequency domain features. The average recognition accuracy of the implemented method for all 21 faults of TEP is 94.08%. To handle multi-time scale data, Gao et al. [23], introduced a feature extraction method based on CNN. The similarity measure is used to eliminate fault-irrelevant variables. The fixed multiple sampling (FMS) method is introduced for converting a feature map with an arbitrary shape into a fixed-length vector to handle different time scales of data. The average FDR for all 21 classes using the multi-time scale model is 98.53%, while the single-time scale model is 94.65%. A multi-model monitoring method is introduced in[41], in which first a CNN model detects some faults quickly and with minimum computational power, then a wavelet-assisted secondary CNN model accurately detects the remaining faults and also filters the noise and redundant information, the FDR of the model for 20 faults of TEP is 93%. A normalized CNN is used in [90] to overcome the problem of gradient explosion and disappearance, moreover the convergence speed of the network is improved. This network consists of the depth-normalized convolution network and improved second-order pooling for feature fusion. The average fault detection accuracy for all faults is 92.75%. It should be noted that this method needs a large amount of calculations. To deal with high dimensional nonlinearity the combination of matrix diagram and multi-scale convolutional neural network (MsCNN) is used in [82]. Time series data is converted into matrix diagrams (2-D images). The average fault detection accuracy for all faults is 88.54%. It should be noted that running this model is very time-consuming. Sparse one-dimensional convolutional neural network (S1-DCNN) [116], is introduced to extract features from complex signals, the method uses a sparsity regularization layer at the output of the convolution network to remove any redundant representations and extract fault-discriminative features. In [73], adaptive multiscale convolutional neural network (AMCNN), two attention modules, adaptive channel attention and adaptive spatial attention, are incorporated to select the most useful features and suppress redundant information. Another novelty in this paper is the proposed optimization method using a weighted loss function consisting of classification loss, regularization loss and triplet loss to improve robustness and generalization. the average fault detection rate for 20 faults of TEP is 93.02%. Dynamic convolutional neural networks can extract temporal correlation of process data and are useful for feature selection, and feature sequence arrangement as in [21], first, the genetic algorithm (GA) is used for feature selection and then CNN is used for fault diagnosis. The FDR for 20 faults of TEP is 89.72%. A 1D-CNN network is used in [60] for FDD. This model exploits the correlation in the data, and the average FDD accuracy for all TEP faults is 93%. Multiblock adaptive convolution kernel neural network (MBCKN) [28] is a feature extraction method for large-scale processes, in which the process is decomposed into several blocks, and the attention mechanism is used for convolutional kernel selection for each block. Finally, an adaptive fusion method fuses the features of different blocks. This model can detect 21 faults of the TEP with an average accuracy of 94.15%. To increase the feature extraction ability and prevent model degradation, a new multi-scale residual jagged dilated convolution neural network (MRJDCNN) model is designed in [15] by incorporating the principle of residual learning into the improved jagged dilated convolution, and then it is combined with the LSTM network delicately. This model could detect all the TEP faults with an average accuracy of 96.59%.
@@ 4.2.1 LSTM and CNN
Yuan et al. [118] proposed a framework for feature extraction and fault diagnosis, where the discrete wavelet transform and CNN are used for feature extraction then the LSTM network reduces the useless information and finally classification is done by using a softmax classifier. The model is validated on a few faults (Fault 1, 2, 4, 8, 10, 12, 13, 14, 17, and 18) of the TEP and the average F1-score is 0.9873. In [79], first, the LSTM network is used for feature extraction, then convolutional kernels extract characteristics of the data. This combination results in the high average fault detection accuracy of 92.06% for the TEP data set. This model can detect faults 3, 9, and 15 which are very difficult to diagnose with a very high accuracy. Bao et al. [6] used a deep convolutional neural network (DCNN) to extract spatial features from raw data and then fuse features into the bidirectional recurrent neural network (BRNN). No additional data preprocessing is needed in this case. The average accuracy for 20 faults of TEP is 91.7%. The combination of the sliding window on integrated 2D image-like data of feature information and time delay information and the CNN-LSTM model is used in [33] for process fault detection. This method is robust against noise and converges faster compared to separate CNN and LSTM networks. To ensure that no feature information is lost, Ren et al. [75] used a CNN to extract multi-level fault features, combine them into a unified feature space, and form hyper features through feature fusion. These hyper features are then fed into a stacked LSTM network for additional feature extraction. This approach effectively captures time series information and addresses the issue of gradient disappearance during long-term training,achieving the accuracy of 95.19% for all the faults. Zhao et al. [131], used an adaptive multiscale CNN and enhanced highway LSTM (ACEL) to tackle the problem of diverse time scales and data features in industrial settings. This method starts by automatically extracting multiscale features under varying operating conditions through a parallel convolutional module. It then applies efficient channel attention to emphasize more important channels while reducing the influence of less relevant ones. A bidirectional LSTM is used to create fine-grained hybrid features, combining contextual and local information learned by the CNN. Additionally, an enhanced highway configuration is employed to strengthen global temporal dependencies. The FDR and F1 are more than 98%.
@@ 4.3 Auto-encoders
Autoencoders are models first introduced in the late 1980s. They consist of two networks, the encoder and the decoder where the encoder maps the input data to a hidden layer and the decoder tries to reconstruct the input data, in general, the process is done unsupervised [78]. They are powerful tools for feature extraction and dimension reduction. The output of the encoder is the feature representation and can be used for further analysis like as the input of a classifier. Different types of autoencoders are shown in Figure 7.
Figure 7: A schematic of different autoencoder variants. Adopted from [84]
@@
Autoencoders have been widely used in condition monitoring and fault detection and diagnosis of industrial plants where the available data is massive, specifically for chemical processes and TEP. One of the first attempts of using autoencoders for fault diagnosis of the TEP is done in [51], a stacked sparse autoencoder (SSAE) is implemented for extracting and learning high-level features and underlying fault patterns which is very useful for detecting incipient faults, moreover time correlations among the samples is also considered, unlike other methods that have poor performance in diagnosing incipient faults, this algorithm could classify faults 3,9,15,21 of the TE which are very hard to detect with the accuracy of 95.75%, 99.25%, 98.75%, 99.5% respectively. Another powerful tool for fault detection especially incipient faults in industrial plants is the Deep LSTM Supervised Autoencoder Neural Network (Deep LSTM-SAE NN) [1]. The assumption that having different models for different faults increases the sensitivity of the model is made in the implementation of this method. The average accuracy for all the classes is 93.23%.
@@
Stacked denoising autoencoder (SDAE) alongside the k-nearest neighbour (KNN) rule is used for automated feature learning in [128], the SDAE automatically extracts crucial features from nonlinear process data. Neural component analysis (NCA) for fault detection is introduced in [129], which is a unified model including a nonlinear encoder and a linear decoder, the network is trained with orthogonal constraints such as those used in PCA to alleviate the overfitting problem in autoencoders. In [17] a variational recurrent autoencoder is used for process fault detection. This method monitors the process in probability space, to handle processes nonlinearity more efficiently. Recurrent neural networks are also used to consider temporal dependency between variables. The negative variational score (NVS) is introduced as the monitoring metric, which considers reconstruction error and similarity between prior and posterior probability distributions. The model is validated on the TEP and the average fault detection rate of all faults expect 3,9,15 is 96.3%. The combination of an autoencoder and LSTM is a powerful tool for fault detection of rare events. In [63], an autoencoder is trained with offline normal data and used as an anomaly detector. The predicted faults will be classified by using an LSTM network. The average fault detection accuracy for the first 20 faults is 91.9%, and by not considering 9 and 15 faults which are very hard to detect, the average accuracy is 96.8%. The time delay of the fault detection of each class is also studied in this work. In [107], a fault detection model based on Mutual Information and Dynamic Stacked Sparse Autoencoder (SSAE) is introduced. The variables with a strong autocorrelation and their lag time of maximum autocorrelation were chosen using time series input data. Then, dynamic augmented data is fed to the SSAE for feature extraction and learning latent information in the dynamic process data. Based on selected process variables the average accuracy of fault detection of all 21 faults is 85.06%. Liu et al. [45], introduced a denoising sparse autoencoder (DSAE) to address the noise interference and nonlinearity in process variables. This method can achieve an accuracy higher than 99% for some faults.

### v3

@@ 3.3 Recurrent Neural Networks
One of the first studies that used LSTM for chemical process fault diagnosis is Zhao et al. (2018), in this case, no specific feature extraction is needed. Applying batch normalization makes the convergence of the network much faster. To improve the diagnosis using LSTM networks, Han et al. Han et al. (2020) used an iterative method to optimize the number of hidden layer nodes. Kang et al. Kang (2020) used RNNs for fault classification. It is shown that RNN can create an even spatial distribution of fault types and classify them accurately. For TEP data, a 4-layer RNN could reach the average classification accuracy of 92.5% for all faults; however, it can not detect fault 15. To overcome this problem, an LSTM network is used, and the average accuracy increased to 95%. LSTM enhances RNNs by incorporating memory cells and gating mechanisms to better capture long-term dependencies. To extract long-term dependencies in raw data and detect abnormal values, a Stacked Long Short-Term Memory network with three layers of LSTM is introduced in Zhang et al. (2020). As the historical information is stored in the memory, the monitoring system can operate in real time. The average FDR for all TEP faults is 99.49%. Chadha et al. Chadha et al. (2020) used bidirectional recurrent neural networks for fault diagnosis, enabling the addition of more information about adjacent points into a training sequence than their unidirectional counterparts. They introduced a novel data preprocessing and restructuring for efficient network training. The average FDR for 21 faults of TEP is 88.14%. To overcome nonlinearity and the high degree of spatio-temporal correlations, and incomplete characterization of the uncertainty in process noise and dynamics, a novel Bayesian recurrent neural network (BRNN) based FDI method is proposed in Sun et al. (2020), where we can have a probabilistic model of input-output mapping. This method can be used for fault propagation path analysis as well. To improve the feature extraction of unidirectional RNNs that only consider positive direction, bidirectional RNN (BiRNN) models are used by Zhang et al. (2019b). Simple RNN unit (SRU), LSTM, and GRU are used in making bidirectional networks with the average FDR of 91.0%, 92.3%, and 92.7%, respectively.
@@
Mirzaei et al. Mirzaei et al. (2022) compared LSTM and GRU networks for FDD of chemical processes by using t-SNE visualization. GRU performed better than LSMT, especially for fault 15, which is very hard to detect. The average accuracies of LSTM and GRU for all faults are 95% and 95.9%, respectively.
@@
Yuan et al. Yuan and Tian (2019a) used GRU to extract the information from correlated sequential data; a moving horizon is used at the input of the GRU to divide the raw data into sequence units. To overcome the covariate shift and improve the network’s learning efficiency, batch normalization is used. BN-based GRU is faster and more accurate. GRU and BN-GRU accuracy is 83.42% and 87.36%, respectively. An end-to-end model fusion feature learning method (MCNN-DBiGRU) is introduced in Zhu et al. (2025), a feature-aligned multi-scale extraction model (MCNN) followed by a deep bidirectional mechanism that extracts time series features in the data, allowing the recurrent neural network to process input features both from past to future and from future to past. The experiments on all the fault samples of TEP achieve the precision of 94.11%.
A summary of recurrent models is shown in Table 3. Recurrent architectures demonstrate strong capability in modeling the temporal dependencies inherent in process signals, making them well-suited for dynamic fault detection tasks. LSTM and GRU models consistently outperform vanilla RNNs due to their gating mechanisms, which mitigate vanishing gradients and enable longer memory retention. GRUs show comparable or slightly better performance in some instances, while offering reduced computational complexity, which is advantageous for real-time applications. However, recurrent models still face limitations in detecting rare faults (e.g., TEP fault 15) unless enhanced with bidirectional processing, attention mechanisms, or combined with convolutional feature extractors.
Table 3: Summary of Recurrent Models. See Table 2 for fault types.
Model Reference Specification Faults Information Average Accuracy
LSTM Yuan and Tian (2019b) Discrete wavelet transform and CNN for feature extraction and softmax classifier 1, 2, 4, 8, 10, 12, 13, 14, 17, and 18 98.73%
BiRNN Zhang et al. (2019b) Bidirectional recurrent neural networks using gated recurrent unit (GRU) All 92.70%
LSTM + CNN Shao et al. (2019) First LSTM and then CNN are used for feature extraction. The detection of faults 3, 9, and 15 are improved. All 92.06%
RNN Kang (2020) Using a 4 layer RNN, however, could not detect fault 15 All 92.50%
LSTM Kang (2020) Better performance compared to RNN speciallly for fault 15 All 95%
LSTM Zhang et al. (2020) Three layers LSTM are stacked All 99.49%
BiRNN Chadha et al. (2020) Bidirectional recurrent neural networks All 88.14%
GRU Mirzaei et al. (2022) The detection of fault 15 is enhanced compared to LSTM All 95.90%
@@ 3.4 Convolutional Neural Network (CNN)
Convolutional Neural Networks (CNNs) were introduced in the late 1980s and initially used to detect patterns in images Le Cun et al. (1989). These powerful feature extraction networks are widely used for fault detection.
@@
One of the first applications of using CNN for TEP fault detection and diagnosis is introduced in Wu and Zhao (2018). In this approach, a deep convolutional neural network (DCNN) with dropout extracts features from time series data. The implemented method achieved an average FDR of 88.2% for 20 TEP faults; however, it heavily relies on historical fault data samples and is less effective for highly complex chemical processes involving thousands of variables. Some modifications to the CNN network are done in Chadha et al. (2019) to increase the generalization ability of the model, improving feature extraction abilities and reducing the overfitting, such as residual connections, inception module, and dilation convolution. With these improvements, the F1 score of hard detectable faults 3,9,13,15 is improved to 0.91. Frequency domain data contains valuable information that is usually neglected in the traditional FDD of chemical processes. In Zhang et al. (2019a), the Fast Fourier Transform (FFT) is used for TEP feature extraction, and the output data is fed into a convolutional neural network, which is used as a binary classifier. The FDR of this model on TEP data, considering all faults except 3, 9, 15, 16, is 96.9%. In CNN, all the features have the same importance in the classification process. Multi-head attention CNN (MACNN) Cui et al. (2020) are introduced to distinguish the importance of different features, where the extracted features from convolution layers are fed into the multi-head attention (MA) layers, finally a softmax classifier is used to classify the different faults, the performance is improved compared to feedforward neural network (FNN) and CNN models. Multiblock temporal convolutional network (MBTCN) for FDD is introduced in He et al. (2021) using the idea of "local extraction and global integration"; it considers the cross-correlation and the temporal-correlation in the multivariate processes’ data. In this method, most related process variables are divided into the same sub-blocks where TCN extracts the features. It achieved 96.2% accuracy in TEP, considering all fault classes but 3,9,15. In Yu et al. (2021), a multichannel one-dimensional convolutional neural network (MC1-DCNN) was employed, using extracted features from the Wavelet transform. The wavelet transform is a powerful tool for extracting time and frequency domain features. The average recognition accuracy of the implemented method for all 21 faults of TEP is 94.08%. To handle multi-time scale data, Gao et al. Gao et al. (2020) introduced a feature extraction method based on CNN. The similarity measure is used to eliminate fault-irrelevant variables. The fixed multiple sampling (FMS) method is introduced for converting a feature map with an arbitrary shape into a fixed-length vector to handle different time scales of data. The average FDR for all 21 classes using the multi-time scale model is 98.53%, while the single-time scale model is 94.65%. A multi-model monitoring method is introduced in Li et al. (2020a), in which first a CNN model detects some faults quickly and with minimum computational power, then a wavelet-assisted secondary CNN model accurately detects the remaining faults and also filters the noise and redundant information. The FDR of the model for 20 faults of TEP is 93%. A normalized CNN is used in Wang et al. (2021) to overcome the problem of gradient explosion and disappearance; moreover, the convergence speed of the network is improved. This network consists of the depth-normalized convolution network and improved second-order pooling for feature fusion. The average fault detection accuracy for all faults is 92.75%. It should be noted that this method needs a large number of calculations. To deal with high-dimensional nonlinearity, the combination of matrix diagram and multi-scale convolutional neural network (MsCNN) is used in Song and Jiang (2022). Time series data is converted into matrix diagrams (2-D images). The average fault detection accuracy for all faults is 88.54%. It should be noted that running this model is very time-consuming. Sparse one-dimensional convolutional neural network (S1-DCNN) Yu et al. (2022b), is introduced to extract features from complex signals, the method uses a sparsity regularization layer at the output of the convolution network to remove any redundant representations and extract fault-discriminative features. In Qin and Zhao (2022a), an adaptive multiscale convolutional neural network (AMCNN), two attention modules, adaptive channel attention and adaptive spatial attention, are incorporated to select the most useful features and suppress redundant information. Another novelty in this paper is the proposed optimization method using a weighted loss function consisting of classification loss, regularization loss, and triplet loss to improve robustness and generalization. The average fault detection rate for 20 faults of TEP is 93.02%. Dynamic convolutional neural networks can extract temporal correlation of process data and are useful for feature selection and feature sequence arrangement, as in Deng et al. (2021). First, the genetic algorithm (GA) is used for feature selection, and then CNN is used for fault diagnosis. The FDR for 20 faults of TEP is 89.72%. A 1D-CNN network is used in Niu and Yang (2022), exploiting the correlation in the data; the average FDD accuracy for all TEP faults is 93%. Multi-block adaptive convolution kernel neural network (MBCKN) Guo et al. (2022) is a feature extraction method for large-scale processes, in which the process is decomposed into several blocks, and the attention mechanism is used for convolutional kernel selection for each block. Finally, an adaptive fusion method fuses the features of different blocks. This model can detect 21 faults of the TEP with an average accuracy of 94.15%. To enhance feature extraction capabilities and prevent model degradation, a novel multi-scale residual jagged dilated convolution neural network (MRJDCNN) model is proposed in Chen et al. (2024). This model incorporates the principle of residual learning into an improved jagged dilated convolution and is then carefully combined with an LSTM network. This model could detect all the TEP faults with an average accuracy of 96.59%. It is also validated on an industrial coking furnace. A global–local attention convolutional neural network (GLACNN) is introduced in Yang et al. (2025) for fault diagnosis in complex multi-unit nonlinear industrial processes, where traditional models struggle to balance intra-unit and inter-unit feature extraction. GLACNN enhances diagnostic performance by combining local features from individual sub-units with global features capturing system-wide interactions, using specially designed one-dimensional convolutional layers. Channel attention modules are employed to assign significance to features based on their relevance, improving interpretability and diagnostic accuracy. Validated on the TEP and penicillin fermentation processes, GLACNN demonstrates superior accuracy compared to baseline models, achieving an average FDR of 99.6% for the first five fault types. However, its performance degrades when encountering unseen fault types common in dynamic industrial settings.
@@ 3.4.1 LSTM and CNN
Yuan et al. Yuan and Tian (2019b) proposed a framework for feature extraction and fault diagnosis, where the discrete wavelet transform and CNN are used for feature extraction, then the LSTM network reduces the useless information, and finally classification is done by using a softmax classifier. The model is validated on a few faults (Fault 1, 2, 4, 8, 10, 12, 13, 14, 17, and 18) of the TEP, and the average F1-score is 0.9873. In Shao et al. (2019), first, the LSTM network is used for feature extraction, then convolutional kernels extract characteristics of the data. This combination results in a high average fault detection accuracy of 92.06% for the TEP data set. This model can detect faults 3, 9, and 15, which are very difficult to diagnose with a very high accuracy. Bao et al. Bao et al. (2022) used a deep convolutional neural network (DCNN) to extract spatial features from raw data and then fuse features into the bidirectional recurrent neural network (BRNN). No additional data preprocessing is needed in this case. The average accuracy for 20 faults of TEP is 91.7%. The combination of the sliding window on integrated 2D image-like data of feature information and time delay information, and the CNN-LSTM model is used in Huang et al. (2022) for process fault detection. This method is robust against noise and converges faster compared to separate CNN and LSTM networks. To ensure that no feature information is lost, Ren et al. Ren et al. (2024) used a CNN to extract multi-level fault features, combine them into a unified feature space, and form hyper features through feature fusion. These hyper features are then fed into a stacked LSTM network for additional feature extraction. This approach effectively captures time series information and addresses the issue of gradient disappearance during long-term training, achieving an accuracy of 95.19% for all the faults. Zhao et al. Zhao et al. (2024) used an adaptive multiscale CNN and enhanced highway LSTM (ACEL) to tackle the problem of diverse time scales and data features in industrial settings. This method starts by automatically extracting multiscale features under varying operating conditions through a parallel convolutional module. It then applies efficient channel attention to emphasize more important channels while reducing the influence of less relevant ones. A bidirectional LSTM is used to create fine-grained hybrid features, combining contextual and local information learned by the CNN. Additionally, an enhanced highway configuration is employed to strengthen global temporal dependencies. The FDR and F1 are more than 98%.
A summary of CNN-based methods is shown in Table 4. Convolutional architectures are powerful in spatial feature extraction from multivariate process data and often outperform simpler feedforward networks. Variants such as MBTCN and MC1-DCNN demonstrate that domain-specific preprocessing (e.g., wavelet transforms, multiblock decomposition) significantly boosts detection rates for complex and multi-timescale processes. Attention-based CNNs (e.g., AMCNN) further improve fault-specific sensitivity by assigning higher weights to critical channels, addressing the equal importance limitation of plain CNNs. The main drawback is computational load, especially for combining CNN with multi-branch attention and multi-resolution networks which can be a barrier to online deployment in high-frequency monitoring scenarios.
Table 4: Summary of CNN Models. See Table 2 for fault types.
Model Reference Specification Faults Information Average Accuracy
DCNN Wu and Zhao (2018) CNN with dropout 20 faults 88.20%
CNN Zhang et al. (2019a) FFT for feature extraction All faults except 3,9,15,16 96.90%
CNN Gao et al. (2020) Handling multi time scale data using fixed multiple sampling method All 98.53%
MBTCN He et al. (2021) Multiblock temporal convolutional network (MBTCN), dividing process variables into subblocks following the idea of local extraction and global integration All faults except 3,9,15 96.20%
MC1-DCNN Yu et al. (2021) 1D CNN and Wavelet transform for feature extraction All 94.08%
Normalized CNN Wang et al. (2021) Improve the stability and speed of the training All 92.75%
MBCKN Guo et al. (2022) Decompose process into several blocks for large-scale process using attention mechanism All 94.15%
CNN + LSTM Ren et al. (2024) CNN is used to extract multi-level fault features, followed by stacked LSTM solving the gradient disappearance problem. All 95.19%
ACEL Zhao et al. (2024) Extracts multiscale features for various operating conditions combining CNN and LSTM All faults except 3,9,15,21 98%
MRJDCNN-LSTM Chen et al. (2024) Improved feature extraction and prevent model degradation All faults 96.59%
GLACNN Yang et al. (2025) Balancing intra-unit and inter-unit feature extraction, assigning significance to features, using channel attention modules, not ideal for unseen faults Faults 1-5 99.6%
@@ 3.5 Auto-encoders
Autoencoders are models first introduced in the late 1980s. They consist of two networks, the encoder and the decoder, where the encoder maps the input data to a hidden layer and the decoder tries to reconstruct the input data. In general, the process is done unsupervised Schmidhuber (2015). They are powerful tools for feature extraction and dimension reduction. The output of the encoder is the feature representation and can be used for further analysis, such as the input of a classifier.
@@
Autoencoders and their variations Ahang et al. (2024) have been widely used in condition monitoring of industrial plants where the available data is massive, specifically for chemical processes. One of the first attempts of using autoencoders for fault diagnosis of the TEP is done in Lv et al. (2017), a stacked sparse autoencoder (SSAE) is implemented for extracting and learning high-level features and underlying fault patterns which is very useful for detecting incipient faults, time correlations among the samples is also considered, unlike other methods that have poor performance in diagnosing incipient faults, this algorithm could classify faults 3,9,15,21 of the TE which are very hard to detect with the accuracy of 95.75%, 99.25%, 98.75%, 99.5% respectively. Another powerful tool for fault detection, especially incipient faults in industrial plants, is the Deep LSTM Supervised Autoencoder Neural Network (Deep LSTM-SAE NN) Agarwal et al. (2022). The assumption that having different models for different faults increases the sensitivity of the model is made in the implementation of this method. The average accuracy for all the classes is 93.23%.
@@
A stacked denoising autoencoder (SDAE) alongside the k-nearest neighbour (KNN) rule is used for automated feature learning in Zhang et al. (2018); the SDAE automatically extracts crucial features from nonlinear process data. Neural component analysis (NCA) for fault detection is introduced in Zhao (2018), which is a unified model including a nonlinear encoder and a linear decoder. The network is trained with orthogonal constraints such as those used in PCA to alleviate the overfitting problem in autoencoders. In Cheng et al. (2019), a variational recurrent autoencoder is used for process fault detection. This method monitors the process in probability space to handle the nonlinearity more efficiently. Recurrent neural networks are also used to consider temporal dependency between variables. The negative variational score (NVS) is introduced as the monitoring metric, which considers reconstruction error and similarity between prior and posterior probability distributions. The model is validated on the TEP and the average fault detection rate of all faults except 3,9,15 is 96.3%. The combination of an autoencoder and an LSTM is a powerful tool for fault detection of rare events. In Park et al. (2019), an autoencoder is trained with offline normal data and used as an anomaly detector. The predicted faults will be classified by using an LSTM network. The average fault detection accuracy for the first 20 faults is 91.9%, and by not considering 9 and 15 faults, which are very hard to detect, the average accuracy is 96.8%. The time delay of the fault detection of each class is also studied in this work. In Yin and Yan (2019), a fault detection model based on Mutual Information and Dynamic Stacked Sparse Autoencoder (SSAE) is introduced. The variables with a strong autocorrelation and their lag time of maximum autocorrelation were chosen using time series input data. Then, dynamic augmented data is fed to the SSAE for feature extraction and learning latent information in the dynamic process data. Based on selected process variables, the average accuracy of fault detection of all 21 faults is 85.06%. Liu et al. Liu et al. (2022) introduced a denoising sparse autoencoder (DSAE) to address the noise interference and nonlinearity in process variables. This method can achieve an accuracy higher than 99% for some faults.